The kinetic energy of an electron accelerated in an x-ray tube is 125 keV. Assuming it is nonrelativistic, what is its wavelength (in m)?
Answer #1

Solution:

Using the formula, K.E = 0.5 * m * v^2

=> 125 * 1.6 * 10^-16 = 0.5 * 9.1 * 10^-31 * v^2

=> v = 2.1 * 10^8 m/s

Now Using the formula,

=> λ = h / mv

=> λ = 6.626 * 10^-34 / (9.1 * 10^-31 * 2.1 * 10^8 )

=> λ = 3.5 * 10^-12 m